OpenStudy (anonymous):

integral of 1/(y(lny)^3) dy

OpenStudy (ash2326):

We have \[\int \frac{1}{y (\ln y)^3} dy\] @ncorrea92 what's the differential of ln x???

OpenStudy (ash2326):

I mean differentiation of ln x?

OpenStudy (anonymous):

1/x

OpenStudy (turingtest):

@ncorrea92 that should suggest a u-substitution to you, try to think what it should be

OpenStudy (anonymous):

would the answer be ln |(lny^3)| + c

OpenStudy (anonymous):

?